# Homeostatic Adaptation of Optimal Population Codes under Metabolic Stress

### v1

@@ Abstract
Information processing in neural populations is inherently constrained by metabolic resource limits and noise properties. Recent data, for example, shows that neurons in mouse visual cortex can go into a “low power mode" in which they maintain firing rate homeostasis while expending less energy. This adaptation leads to increased neuronal noise and tuning curve flattening in response to metabolic stress. These dynamics are not described by existing mathematical models of optimal neural codes. We have developed a theoretical population coding framework that captures this behavior using two surprisingly simple constraints: an approximation of firing rate homeostasis and an energy limit tied to noise levels via biophysical simulation. A key feature of our contribution is an energy budget model directly connecting adenosine triphosphate (ATP) use in cells to a fully explainable mathematical framework that generalizes existing optimal population codes. Specifically, our simulation provides an energy-dependent dispersed Poisson noise model, based on the assumption that the cell will follow an optimal decay path to produce the least-noisy spike rate that is possible at a given cellular energy budget. Each state along this optimal path is associated with properties (resting potential and leak conductance) which can be measured in electrophysiology experiments and have been shown to change under prolonged caloric deprivation. We analytically derive the optimal coding strategy for neurons under varying energy budgets and coding goals, and show that our method uniquely captures how populations of tuning curves adapt while maintaining homeostasis, as has been observed empirically.
@@ 1 Introduction
Animals have limited access to calories, which constrains their brains’ ability to expend energy to represent information. Additionally, energy budgets are inconsistent over the lifetime of an organism, which must adapt in times of scarcity while maintaining as much functionality as possible. Padamsey et al. have recently shown that under long-term calorie deprivation, mice lose the ability to discriminate fine-grained visual details while keeping coarser information, and that they do so with systematic changes in biophysical cell properties that result in a flattening [^1] While the change is described in [22] in terms of widening, firing rate homeostasis requires that peak firing rates drop in tandem with this change. This difference is significant: a Gabor that widens changes its preferred frequency, but a Gabor that flattens does not. of neuronal tuning curves [22]. Perhaps surprisingly, this energy-saving strategy does not result in a significant change in spike counts, demonstrating the importance of firing rate homeostasis even when energy budgets are tight.
The efficient coding hypothesis [1] asserts that populations of sensory neurons are optimized to maximize information about the environment subject to resource constraints. While the impact of coding objectives [5, 6, 18, 30] and noise properties [15, 18, 3, 30] on optimal population codes have been studied in depth, the effect of energy constraints and in particular their variation over time remains underexplored. Simple models of energy constraints (e.g., on mean or maximum firing rates) are analytically convenient, but they fail to capture the complex trade-offs determining energetic costs in real cells, such as subthreshold activity and biophysical adaptations [21]. We will show that these simplified constraints lead to inaccurate predictions of neurons’ adaptation to metabolic stress (i.e., reduction of energy budget). Specifically, our model predicts tuning curve flattening closely in line with [22], while previous population coding models capture either the shortening or widening behavior of energy-limited tuning curves, but not both.
@@
  • We introduce two mathematically-tractable constraints that more accurately characterize real neurons’ response to metabolic stress.
@@
Our paper is organized as follows: Sect. 2 provides context from the literature, Sect. 3 presents our novel constraints and shows how our analytical model makes new predictions while generalizing previous methods, Sect. 4 confirms the validity of these constraints with biophysical simulations that set values of parameters introduced in Sect. 3, and Sect. 5 demonstrates that our simulation-grounded analytical model uniquely predicts the changes seen in metabolically-stressed mouse data.
@@ 2 Related work
Since the foundational work by Barlow [1], the efficient coding hypothesis has been extensively evaluated across a variety of contexts, leading to the development of numerous information-theoretic models with strong empirical support from both neural and perceptual data [7, 18, 27, 24, 36, 14]. A variety of coding goals have been analyzed, from entropy maximization [14] to mutual information maximization [2, 33, 31, 18, 5, 6] to $L_{p}$ norm error minimization [30, 5, 6, 18].
Optimal coding models are often built on Fisher information (FI) [30, 5, 6, 18, 31], which describes the local sensitivity of a population response around a stimulus value. One reason for this is that FI can be estimated more easily from measurements of tuning curves than entropy-based information metrics. Another is that FI optimization is closely related to standard coding objectives: maximizing $\log$ FI approximates mutual information maximization [2, 33], while maximizing power functions of FI minimizes $L_{p}$ loss in the asymptotic limit [29]. Coding goals of particular interest here are the "infomax" (mutual information maximization, through $\log$ FI maximization) and "discrimax" (discriminability maximization, through $FI^{-1}$ maximization, implying $L_{2}$ error minimization by the Cramér Rao bound) objectives.
Optimal neuronal populations can be predicted directly from FI-based objectives, provided that significant simplifying assumptions are imposed on tuning curve shapes. For example, [32] assumes homogeneous tuning widths or amplitudes across the population. Usefully, Ganguli and Simoncelli [5] introduced a tiling assumption that enables a more compact representation of population codes through two continuous functions: gain and density, denoted by $g(\cdot)$ and $d(\cdot)$, respectively. These functions are then used to modulate a base shape (e.g., Gaussian or sigmoid) to construct heterogeneous population codes. Wang et al. [29] also adopt the tiling assumption, with a fixed gain function.
@@
Constraints shape optimal codes just as strongly as coding objectives. Constraints in prior literature can typically be categorized into three types: energy, coding capacity, and population size. Energy-based constraints, such as those based on mean firing rate [5, 35, 23] or maximum firing rate [30, 31, 14], reflect the energetic costs of neuronal firing. Coding capacity, expressed as the sum of the square root of FI, has been employed in [33, 29] and generalized to other power laws in [20]. The population size constraint [5, 35] reflects the finite number of neurons available during neural processing.
The joint influence of coding objectives and constraints on shaping optimal population codes has also been explored in recent work. Morais and Pillow [20] show how direct trade-offs between coding objectives and FI-based constraints can obscure the relationship between these factors. Rast and Drugowitsch [23] describe an adaptation procedure to jointly determine objectives and constraints from neural data, but their model does not allow changes in co-firing patterns like those caused by tuning curve widening. Most similar to our work, [31] analyzes ON-OFF population codes under metabolic stress by modeling both energy costs and noise levels as power laws of firing rate. However, their energy cost model is inconsistent with [22] and they do not connect the two terms through a noise-energy trade-off. None of these constraints connect directly to ATP use, due to the biophysical complexity of real neurons.
@@
ATP is the energy currency of cells, yet most models of energy consumption in neurons use the simplifying assumption that spikes have a fixed cost and dominate the neuron’s ATP usage. In fact, spikes are not typically the largest part of a neuron’s ATP budget, costing only 22% of energy use [10]. Reversal of ion fluxes from postsynaptic receptors (50%) and reversal of leak sodium entry via the Na+/K+ exchanger (20%) together dominate ATP usage in neurons [10]. We use a biophysical model of a neuron that estimates ATP usage from all three of these sources, creating a more realistic energy constraint than spike rate alone.
In fact, one of the key features of our neuron simulation is that it estimates different energy rates while maintaining a fixed spike rate. This result is consistent with [22] and relies on systematic changes in cell parameters observed in metabolically stressed mouse cortex. Homeostatic maintenance of a fixed firing rate is a commonly observed feature in neural circuits, often involving the balance of excitation and inhibition [16].
@@ 3 Analytical Model
Based on the neuron simulation described in Sec.˜4, energy-varying noise can be expressed as a dispersed Poisson distribution, with a dispersion factor $\eta_{\kappa}(E)$} that captures variability in neural responses under different activity levels $\kappa$ and energy budgets $E$:
@@ 3.1 Approximation with tiling
To approximate the summation in eq.˜3 in terms of $g(s)$ and $d(s)$, we adopt the same tiling property as in [5, 29]. Under this assumption, the Fisher information is approximated as:
@@
$p(s)\frac{g(s)}{d(s)}=R(s).$ (5)
This approximate homeostasis constraint can be visualized as a rectangular (height $g(s_{n})$, width $1/d(s_{n})$) approximation of each Gaussian tuning curve and a locally constant $p(s)$. Detailed derivations of eqs. 4 and 5 are provided in Apps.˜C, B and A.
Combining the approximations in eq.˜4 and eq.˜5 gives the following population optimization:
$\operatorname*{argmax}_{g(\cdot),d(\cdot)}\int p(s)f\left(\eta_{\kappa}(E)^{-1}g(s)d(s)^{2}\right)\mathrm{d}s,\quad\text{s.t.}\quad p(s)\frac{g(s)}{d(s)}=R(s).$ (6)
@@
This model is consistent with [22], which shows that energy savings come from reducing synaptic conductance (analogous to gain) rather than other cell parameters that widen the tuning curve (analogous to density), which we interpret as compensation to maintain homeostasis. We will relate the energy budget $E$ to physical energy expenditure and fix the value of $\alpha$ for real neurons in Sec.˜4.
@@ 3.2 Analytical solution
We analytically solve the proposed optimization framework for various objective functions (see Tab.˜1) by eliminating $d(s)$ via the approximate homeostasis constraint and applying the Lagrangian method. Here, we outline the key steps in the derivation for the infomax case, with other objective functions and full details included in App.˜D. To solve the optimization problem defined by eqs.˜6 and 7, we eliminate $d(s)$ by substituting from the approximate homeostasis constraint in eq.˜5, yielding the following formulation:
@@
Setting $\partial_{g}\mathcal{L}=0$ and using eq.˜7 provides the optimal gain, with density following from eq.˜5:
@@
The resulting FI is obtained with eq.˜4 and the discriminability bound $\delta_{\min}$ from the Cramér Rao bound, summarized in Tab.˜1, along with the analytical solutions corresponding to $L_{p}$ error objective functions.
Table 1: Analytical solutions in our framework. For a comparison to existing methods, see App.˜K.
@@
Density (tuning width)-1 $d(s)$ $E^{\frac{1}{\alpha}}R(s)^{-1}p(s)$ $\propto E^{\frac{1}{\alpha}}R(s)^{\frac{-\alpha-1}{\alpha+3}}p(s)^{\frac{\alpha+1}{\alpha+3}}$ $\propto E^{\frac{1}{\alpha}}R(s)^{\frac{\alpha-\beta}{3\beta-\alpha}}p(s)^{\frac{\beta-\alpha}{3\beta-\alpha}}$
@@
Fisher information $FI(s)$ $\propto\frac{E^{\frac{3}{\alpha}}p(s)^{2}}{\eta_{\kappa}(E)R(s)^{2}}$ $\propto\frac{E^{\frac{3}{\alpha}}p(s)^{\frac{2\alpha}{\alpha+3}}}{\eta_{\kappa}(E)R(s)^{\frac{2\alpha}{\alpha+3}}}$ $\propto\frac{E^{\frac{3}{\alpha}}p(s)^{\frac{-2\alpha}{3\beta-\alpha}}}{\eta_{\kappa}(E)R(s)^{\frac{2\alpha}{\alpha-3\beta}}}$
Discriminability bound $\delta_{\min}(s)$ $\propto E^{\frac{-3}{2\alpha}}p(s)^{-1}$ $\propto E^{\frac{-3}{2\alpha}}p(s)^{\frac{-\alpha}{\alpha+3}}$ $\propto E^{\frac{-3}{2\alpha}}p(s)^{\frac{\alpha}{3\beta-\alpha}}$
@@ 3.3 Relation to existing models
The approximate homeostasis and energy constraints introduced in eqs.˜5 and 7 are not only grounded in biophysiological evidence but also generalize commonly used constraints in existing models (see Tab.˜2). Here, we first demonstrate how our constraints relate to those in [5], followed by a discussion of the connection to [29]. It is evident that our energy constraint reduces to the mean firing rate (FR) constraint used in [5] when setting $\alpha=1$ and redefining $E$ as the mean firing rate denoted as $R$. The population size constraint can also be recovered from our constraints as follows:
$E\overset{\alpha:=1}{=}\int p(s)g(s)\,\mathrm{d}s\overset{\text{approx. homeo.}}{=}\int R(s)d(s)\,\mathrm{d}s\overset{R(s):=\frac{E}{N}}{=}\frac{E}{N}\int d(s)\,\mathrm{d}s\quad\Rightarrow\quad\int d(s)\,\mathrm{d}s=N.$ (12)
By jointly enforcing the conditions that yield the mean FR and population size constraints, we obtain $R(s)=R/N$, indicating that the population mean firing rate $R$ is evenly distributed across neurons. It is important to note that $R(s)$ denotes the firing rate at the level of stimulus values, while $R$ represents the overall firing rate in the framework of [5]. This leads us to observe that approximate homeostasis, though not an explicit requirement of [5], emerges as a signature of optimality in their model (see App.˜E for proof of this proposition).
To reduce our constraint to the constraint of total FI states in [29], we now set $\alpha=3/2$:
@@
where $C\in\mathbb{R}$ is the FI coding capacity. Under the assumption $\alpha=3/2$, the above reduction implies that our energy constraint can be interpreted as a weighted version of a coding resource constraint, where the weighting function corresponds to the firing rate $R(s)$. The generalization of our model to the existing frameworks is summarized and illustrated in Fig.˜1.
@@
Ganguli & Simoncelli [5] mean FR population size $\int p(s)g(s)\mathrm{d}s=R$ $\int d(s)\mathrm{d}s=N$ Poisson $R\downarrow$, shortens ($N$ fixed)
Wang et al. [30] max FR total FI states $g(s)\propto 1$ $\int\sqrt{g(s)d(s)^{2}}\mathrm{d}s=C$ Poisson or Gaussian $C\downarrow$, widens ($g$ fixed)
@@
Figure 1: A general solution. Our analytical framework generalizes previous work [5, 30] and predicts tuning curve flattening with a biophysical grounding in neuron simulation.

### v2

@@ Abstract
Information processing in neural populations is inherently constrained by metabolic resource limits and noise properties. Recent data, for example, shows that neurons in mouse visual cortex can go into a “low power mode” in which they maintain firing rate homeostasis while expending less energy. This adaptation leads to increased neuronal noise and tuning curve flattening in response to metabolic stress. These dynamics are not described by existing mathematical models of optimal neural codes. We have developed a theoretical population coding framework that captures this behavior using two surprisingly simple constraints: an approximation of firing rate homeostasis and an energy limit tied to noise levels via biophysical simulation. A key feature of our contribution is an energy budget model directly connecting adenosine triphosphate (ATP) use in cells to a fully explainable mathematical framework that generalizes existing optimal population codes. Specifically, our simulation provides an energy-dependent dispersed Poisson noise model, based on the assumption that the cell will follow an optimal decay path to produce the least-noisy spike rate that is possible at a given cellular energy budget. Each state along this optimal path is associated with properties (resting potential and leak conductance) which can be measured in electrophysiology experiments and have been shown to change under prolonged caloric deprivation. We analytically derive the optimal coding strategy for neurons under varying energy budgets and coding goals, and show that our method uniquely captures how populations of tuning curves adapt while maintaining homeostasis, as has been observed empirically.
@@ 1 Introduction
Animals have limited access to calories, which constrains their brains’ ability to expend energy to represent information. Additionally, energy budgets are inconsistent over the lifetime of an organism, which must adapt in times of scarcity while maintaining as much functionality as possible. Padamsey et al. have recently shown that under long-term calorie deprivation, mice lose the ability to discriminate fine-grained visual details while keeping coarser information, and that they do so with systematic changes in biophysical cell properties that result in a flattening of neuronal tuning curves (padamsey2022neocortex). Perhaps surprisingly, this energy-saving strategy does not result in a significant change in spike counts, demonstrating the importance of firing rate homeostasis even when energy budgets are tight.
This energy-saving strategy motivates us to develop an analytical and simulation-grounded framework that describes how metabolically-stressed neurons can optimally change to produce noisier, cheaper spike trains rather than reducing their firing rates, consistent with Padamsey et al.’s empirical observations. Firing rate homeostasis under metabolic stress, along with direct but tractable accounting for ATP use, is the key to correctly modeling how real tuning curves change. The standard analytical approaches of limiting spike rates (ganguli2010implicit; wang2016efficient) can predict optimal populations under a single metabolic state, but fails to capture how codes adapt to changing conditions.
The efficient coding hypothesis (barlow1961possible) asserts that populations of sensory neurons are optimized to maximize information about the environment subject to resource constraints. While the impact of coding objectives (ganguli2010implicit; ganguli2014efficient; manning2024transformations; wang2016efficient) and noise properties (laughlin1998metabolic; manning2024transformations; ecker2011effect; wang2016efficient) on optimal population codes have been studied in depth, the effect of energy constraints and in particular their variation over time remains underexplored. Simple models of energy constraints (e.g., on mean firing rate in ganguli2010implicit; yerxa2020efficient; rast2020adaptation or maximum firing rate in wang2016efficient; wang2016metab; laughlin1981simple) are analytically convenient, but they fail to capture the complex trade-offs determining energetic costs in real cells, such as subthreshold activity and biophysical adaptations (padamsey2023paying). We will show that these simplified constraints lead to inaccurate predictions of neurons’ adaptation to metabolic stress (i.e., reduction of energy budget). Specifically, our model predicts tuning curve flattening closely in line with padamsey2022neocortex, while previous population coding models capture either the shortening (ganguli2010implicit) or widening (wang2016efficient) behavior of energy-limited tuning curves, but not both.
@@
  • We introduce two mathematically-tractable constraints that are the first to accurately describe real neurons’ recently-characterized response to metabolic stress.
@@ 2 Related work
Since the foundational work by barlow1961possible, the efficient coding hypothesis has been extensively evaluated across a variety of contexts, leading to the development of numerous information-theoretic models with strong empirical support from both neural and perceptual data (ganguli2016neural; manning2024transformations; Vacher2023PerceptualSP; risau2020colour; zhaoping2011human; laughlin1981simple). A variety of coding goals have been analyzed, from entropy maximization (laughlin1981simple) to mutual information maximization (brunel1998mutual; wei2016mutual; wang2016metab; manning2024transformations; ganguli2010implicit; ganguli2014efficient) to $L_{p}$ norm error minimization (wang2016efficient; ganguli2010implicit; ganguli2014efficient; manning2024transformations).
Optimal coding models are often built on Fisher information (FI) (wang2016efficient; ganguli2010implicit; ganguli2014efficient; manning2024transformations; wang2016metab), which describes the local sensitivity of a population response around a stimulus value. One reason for this is that FI can be estimated more easily from measurements of tuning curves than entropy-based information metrics. Another is that FI optimization is closely related to standard coding objectives: maximizing $\log$ FI approximates mutual information maximization (brunel1998mutual; wei2016mutual), while maximizing power functions of FI minimizes $L_{p}$ loss in the asymptotic limit (wang2012optimal). Coding goals of particular interest here are the ”infomax” (mutual information maximization, through $\log$ FI maximization) and ”discrimax” (discriminability maximization, through $FI^{-1}$ maximization, implying $L_{2}$ error minimization by the Cramér Rao bound) objectives.
Optimal neuronal populations can be predicted directly from FI-based objectives, provided that significant simplifying assumptions are imposed on tuning curve shapes. For example, wei2015bayesian assumes homogeneous tuning widths or amplitudes across the population. Usefully, ganguli2010implicit introduced a tiling assumption that enables a more compact representation of population codes through two continuous functions: gain and density, denoted by $g(\cdot)$ and $d(\cdot)$, respectively. These functions are then used to modulate a base shape (e.g., Gaussian or sigmoid) to construct heterogeneous population codes. wang2012optimal also adopt the tiling assumption, with a fixed gain function.
@@
Constraints shape optimal codes just as strongly as coding objectives. Constraints in prior literature can typically be categorized into three types: energy, coding capacity, and population size. Energy-based constraints, such as those based on mean firing rate (ganguli2010implicit; yerxa2020efficient; rast2020adaptation) or maximum firing rate (wang2016efficient; wang2016metab; laughlin1981simple), reflect the energetic costs of neuronal firing. Coding capacity, expressed as the sum of the square root of FI, has been employed in wei2016mutual; wang2012optimal and generalized to other power laws in morais2018power. The population size constraint ganguli2010implicit; yerxa2020efficient reflects the finite number of neurons available during neural processing.
The joint influence of coding objectives and constraints on shaping optimal population codes has also been explored in recent work. morais2018power show how direct trade-offs between coding objectives and FI-based constraints can obscure the relationship between these factors. rast2020adaptation describe an adaptation procedure to jointly determine objectives and constraints from neural data, but their model does not allow changes in co-firing patterns like those caused by tuning curve widening. Most similar to our work, wang2016metab analyzes ON-OFF population codes under metabolic stress by modeling both energy costs and noise levels as power laws of firing rate. However, their energy cost model is inconsistent with padamsey2022neocortex and they do not connect the two terms through a noise-energy trade-off. None of these constraints connect directly to ATP use, due to the biophysical complexity of real neurons.
@@
ATP is the energy currency of cells, yet most models of energy consumption in neurons use the simplifying assumption that spikes have a fixed cost and dominate the neuron’s ATP usage. In fact, spikes are not typically the largest part of a neuron’s ATP budget, costing only 22% of energy use (harris2012synaptic). Reversal of ion fluxes from postsynaptic receptors (50%) and reversal of leak sodium entry via the Na+/K+ exchanger (20%) together dominate ATP usage in neurons (harris2012synaptic). We use a biophysical model of a neuron that estimates ATP usage from all three of these sources, creating a more realistic energy constraint than spike rate alone.
In fact, one of the key features of our neuron simulation is that it estimates different energy rates while maintaining a fixed spike rate. This result is consistent with padamsey2022neocortex and relies on systematic changes in cell parameters observed in metabolically stressed mouse cortex. Homeostatic maintenance of a fixed firing rate is a commonly observed feature in neural circuits, often involving the balance of excitation and inhibition (Liang2024).
It is worth noting that the approximating a neuronal energy cost through firing rate is commonly adopted in recent literature (tavoni2025convergence; koren2022biologically; koren2025efficient; gutierrez2019population) for analyzing changes in intrinsic cell-level and/or network level properties. Although these works use firing rate as an energy proxy same as in ganguli2010implicit; yerxa2020efficient; rast2020adaptation; wang2016efficient; wang2016metab; laughlin1981simple, they do not consider the effect of metabolic stress on neurons, as reported in padamsey2022neocortex. In fact, these models cannot reiterate a central finding of this recent work: that spike rates remain the same as energy use is reduced.
Recently-discovered neural mechanism.
padamsey2022neocortex show that neurons change their biophysical state so that they can maintain their mean firing rates while expending less energy. Primarily, they modify electrical properties to reduce the current associated with receiving input spikes, because reversing this current costs significant ATP through ion pumps. Essentially, by sitting closer to the spiking threshold and receiving a smaller signal from each incoming spike, the cell uses less energy for signal-carrying spikes but also becomes readier to fire in response to stochastic inputs and system noise, which makes them more vulnerable to false firings. Because these false firings are strictly positive, they raise the zero values at the edges of their tuning curves (widening), which requires a compensatory drop in peak values (flattening) to maintain homeostasis.
@@ 3 Analytical Model
Based on the neuron simulation described in Sec. 4, energy-varying noise can be expressed as a dispersed Poisson distribution, with a dispersion factor $\eta_{\kappa}(E)$} that captures variability in neural responses under different activity levels $\kappa$ and energy budgets $E$:
@@ 3.1 Approximation with tiling
To approximate the summation in eq. 3 in terms of $g(s)$ and $d(s)$, we adopt the same tiling property as in ganguli2010implicit; wang2012optimal. Under this assumption, the Fisher information is approximated as:
@@
$p(s)g(s)=R(s)d(s).$ (5)
This approximate homeostasis constraint can be visualized as a rectangular (height $g(s_{n})$, width $1/d(s_{n})$) approximation of each Gaussian tuning curve and a locally constant $p(s)$. Detailed derivations of eqs. 4 and 5 are provided in Apps. C, B and A.
Combining the approximations in eq. 4 and eq. 5 gives the following population optimization:
$\operatorname*{argmax}_{g(\cdot),d(\cdot)}\int p(s)f\left(\eta_{\kappa}(E)^{-1}g(s)d(s)^{2}\right)\mathrm{d}s,\quad\text{s.t.}\quad p(s)g(s)=R(s)d(s).$ (6)
@@
This model is consistent with padamsey2022neocortex, which shows that energy savings come from reducing synaptic conductance (analogous to gain) rather than other cell parameters that widen the tuning curve (analogous to density), which we interpret as compensation to maintain homeostasis. We will relate the energy budget $E$ to physical energy expenditure and fix the value of $\alpha$ for real neurons in Sec. 4.
@@ 3.2 Analytical solution
We analytically solve the proposed optimization framework for various objective functions (see Tab. 1) by eliminating $d(s)$ via the approximate homeostasis constraint and applying the Lagrangian method. Here, we outline the key steps in the derivation for the infomax case, with other objective functions and full details included in App. D. To solve the optimization problem defined by eqs. 6 and 7, we eliminate $d(s)$ by substituting from the approximate homeostasis constraint in eq. 5, yielding the following formulation:
@@
Setting $\partial_{g}\mathcal{L}=0$ and using eq. 7 provides the optimal gain, with density following from eq. 5:
@@
The resulting FI is obtained with eq. 4 and the discriminability bound from the Cramér Rao bound, summarized in Tab. 1, along with the analytical solutions corresponding to $L_{p}$ error objective functions. It is interesting to point out that our model predicts that both density (i.e., $1/\text{width}$) and gain scale on the order of $E^{1/\alpha}$, while the Fisher information scales on the order of $\eta_{\kappa}(E)^{-1}\,E^{3/\alpha}$. Additionally, we show in App. N that the analytical solution in eqs. 10 and 11 are robust to differential correlations, which are the dominant violation of the independent noise assumption in real neurons (moreno2014information). This robustness extends to significant changes in noise correlations across metabolic states.
Table 1: Analytical solutions in our framework. For comparison to other methods, see App. K.
@@
Density $d(s)$ $E^{\frac{1}{\alpha}}R(s)^{-1}p(s)$ $\propto E^{\frac{1}{\alpha}}R(s)^{\frac{-\alpha-1}{\alpha+3}}p(s)^{\frac{\alpha+1}{\alpha+3}}$ $\propto E^{\frac{1}{\alpha}}R(s)^{\frac{\alpha-\beta}{3\beta-\alpha}}p(s)^{\frac{\beta-\alpha}{3\beta-\alpha}}$
@@
Fisher info $FI(s)$ $\propto\frac{E^{\frac{3}{\alpha}}p(s)^{2}}{\eta_{\kappa}(E)R(s)^{2}}$ $\propto\frac{E^{\frac{3}{\alpha}}p(s)^{\frac{2\alpha}{\alpha+3}}}{\eta_{\kappa}(E)R(s)^{\frac{2\alpha}{\alpha+3}}}$ $\propto\frac{E^{\frac{3}{\alpha}}p(s)^{\frac{-2\alpha}{3\beta-\alpha}}}{\eta_{\kappa}(E)R(s)^{\frac{2\alpha}{\alpha-3\beta}}}$
Disc. bound $\propto E^{\frac{-3}{2\alpha}}p(s)^{-1}$ $\propto E^{\frac{-3}{2\alpha}}p(s)^{\frac{-\alpha}{\alpha+3}}$ $\propto E^{\frac{-3}{2\alpha}}p(s)^{\frac{\alpha}{3\beta-\alpha}}$
@@ 3.3 Relation to existing models
The approximate homeostasis and energy constraints introduced in eqs. 5 and 7 are not only grounded in biophysiological evidence but also generalize commonly used constraints in existing models (see Tab. 2). Here, we first demonstrate how our constraints relate to those in ganguli2010implicit, followed by a discussion of the connection to wang2012optimal. It is evident that our energy constraint reduces to the mean firing rate (FR) constraint used in ganguli2010implicit when setting $\alpha=1$ and redefining $E$ as the mean firing rate denoted as $R$. The population size constraint can also be recovered from our constraints as follows:
$E\overset{\alpha:=1}{=}\int p(s)g(s)\,\mathrm{d}s\overset{\text{approx. homeo.}}{=}\int R(s)d(s)\,\mathrm{d}s\overset{R(s):=\frac{E}{N}}{=}\frac{E}{N}\int d(s)\,\mathrm{d}s\Rightarrow\int d(s)\,\mathrm{d}s=N.$ (12)
By jointly enforcing the conditions that yield the mean FR and population size constraints, we obtain $R(s)=R/N$, indicating that the population mean firing rate $R$ is evenly distributed across neurons. It is important to note that ganguli2010implicit uses the parameter $R$ to denote the total expected firing rate of a neuron population. The parameter $R_{n}$ in eq. 3 represents the expected firing rate of the $n^{\mathrm{th}}$ neuron in the population. Meanwhile, $R(s)$ can be interpreted as the expected firing rate of a neuron whose preferred stimulus is $s$, and can be regarded as a “continuous” version of $R_{n}$. The derivation in eq. 12 leads us to observe that approximate homeostasis, though not an explicit requirement of ganguli2010implicit, emerges as a signature of optimality in their model (see App. E for proof of this proposition).
To reduce our constraint to the constraint of total FI states in wang2012optimal, we now set $\alpha=3/2$:
@@
where $C\in\mathbb{R}$ is the FI coding capacity. Under the assumption $\alpha=3/2$, the above reduction implies that our energy constraint can be interpreted as a weighted version of a coding resource constraint, where the weighting function corresponds to the firing rate $R(s)$. The generalization of our model to the existing frameworks is summarized and illustrated in Fig. 1.
In summary, we show that our general case can be reduced to previous models in the literature. Specifically, by setting $\alpha=1$ and using firing rate as an energy proxy reduces our constraint to that of ganguli2010implicit. Furthermore, when $\alpha=3/2$ and the FI coding capacity is incorporated, our energy constraint reduces to the coding capacity constraint proposed in wang2012optimal. In the following section, we will describe the biophysical parameters that reduce the general case to a true ATP-based energy constraint.
@@
Ganguli & Simoncelli (2010) mean FR population size $\int p(s)g(s)\mathrm{d}s=R$ $\int d(s)\mathrm{d}s=N$ Poisson $R\downarrow$, shortens ($N$ fixed)
wang2016efficient max FR total FI states $g(s)\propto 1$ $\int\sqrt{g(s)d(s)^{2}}\mathrm{d}s=C$ Poisson or Gaussian $C\downarrow$, widens ($g$ fixed)
@@
Figure 1: A general solution. Our analytical framework generalizes previous work (ganguli2010implicit; wang2016efficient) and predicts tuning curve flattening with a biophysical grounding in neuron simulation.